A car of 700 kg moving with a velocity of 10 m/s. A motor bike of 120 kg




also moving with same speed in the same direction lying behind 10 meter




from the car. What is there collision velocity?(

1
Expert's answer
2022-03-13T18:48:23-0400

"u(m_1+m_2)=vm_1+vm_2,"

"u=\frac{vm_1+vm_2}{m_1+m_2}=10~\frac ms."